1999 TEX Users Group Election

Election Notice

The terms of the TUG President and of 10 members of the Board of Directors expire as of the 1999 Annual Board Meeting, which will take place in conjunction with the 20th Annual TUG Meeting to be held 15–19 August 1999 in Vancouver, Canada.

The current President, Mimi Jett, has stated her intention to stand for re-election. The directors whose terms expire in 1999 are Kristoffer Rose, Don DeLand, Barbara Beeton, Karl Berry, Kaja Christiansen, Susan DeMeritt, Judy Johnson, Ross Moore, Cameron Smith and Philip Taylor; directors named in italic are expected to stand for election to another term. Continuing directors, with terms ending in 2001, are Arthur Ogawa, Patricia Monohon, and Petr Sojka.

The election to choose the new President and Board members will be held in Spring of 1999. Nominations for these openings are now being invited.

The Bylaws provide that “Any member may be nominated for election to the office of TUG President/to the Board by submitting a nomination petition in accordance with the TUG Election Procedures. Election … shall be by written mail ballot of the entire membership, carried out in accordance with those same Procedures.”

The name of any member may be placed in nomination for election to one of the open offices by submission of a petition, signed by two other members in good standing, to the TUG office at least two weeks (14 days) prior to the mailing of ballots. (A candidate’s membership dues for 1999 will be expected to be paid by the nomination deadline.) A nomination form follows this announcement; forms may also be obtained from the TUG office, and electronically via the TUG Web pages at

http://www.tug.org/1999elect.html.

Along with a nomination form, each candidate is asked to supply a passport-size photograph, a short biography, and a statement of intent to be included with the ballot; the biography and statement of intent together may not exceed 400 words. The deadline for receipt at the TUG office of nomination forms and ballot information is 15 March 1999 at 1700 PST (–8).

Ballots will be mailed to all members about 30 days after the close of nominations. Marked ballots must be postmarked no more than six (6) weeks following the mailing; the exact dates will be noted on the ballots.

Ballots will be counted by a disinterested party not part of the TUG organization. The results of the election should be available early in June, and will be announced in a future issue of TUGboat as well as through various TEX-related electronic lists.

It is the intent of TUG that all wishing to participate in TUG elections do so. If you experience difficulty in submitting your nomination, do not receive your ballot in a timely way, or experience other problems, please contact the TUG office or mailto:election@tug.org.

Barbara Beeton

for the Elections Committee

Nomination Form

Only TUG members whose dues have been paid for 1999 will be eligible to participate in the election. The signatures of two (2) members in good standing at the time they sign the nomination form are required in addition to that of the nominee. Type or print names clearly; both supporting signatures are required. Please use the name by which you are known to TUG; all names must be identified from the TUG membership records before being accepted as valid. Please provide an email address if you wish an electronic confirmation of receipt of your nomination or supporting signature.

The undersigned TUG members propose the nomination of:

Print name of Nominee

signature E-mail or FAX

date

for the position of (check one):

TUG President

Member of the TUG Board of Directors for a term beginning with the 1999 Annual Meeting, August 1999.

Members supporting this nomination:

1.  _{please print name}_{signature}  _{E-mail or FAX}_{date}

2.  _{please print name}_{signature}  _{E-mail or FAX}_{date}

Return this nomination form to the TUG office (FAXed forms will be accepted). Nomination forms and all required supplementary material (photograph, biography and personal statement for inclusion on the ballot) must be received in the TUG office no later than 15 March 1999.^{1} It is the responsibility of the candidate to ensure that this deadline is met; candidates are advised to check with the TUG office that their nomination forms are in order.

nomination form

photograph

biography/personal statement

TEX Users Group FAX: +1 503 223-3960 Nominations for 1999 Election P.O. Box 2311 Portland, OR 97208–2311 USA

1Supplementary material may be sent separately from the form, and supporting signatures need not all appear on one form.